OpenStudy (anonymous):

The legs of a right triangle are 10 centimeters and 24 centimeters long. What is the length of the hypotenuse? 22 centimeters 26 30 centimeters 34 centimeters

OpenStudy (astrophysics):

 use pythagorean theorem \[x^2+y^2 = r^2 \implies r = \sqrt{x^2+y^2}\]